IN A PEN OF GOATS AND CHICKENS, THERE ARE 40 HEADS AND 130 FEET. HOW MANY GOATS AND CHICKENS ARE IN THE PEN??


Answer by nerdybill(5412)

IN A PEN OF GOATS AND CHICKENS, THERE ARE 40 HEADS AND 130 FEET. HOW MANY GOATS AND CHICKENS ARE IN THE PEN??
.
Let p = number of pigs
then
40-p = number of goats
.
4p + 2(40-p) = 130
2p + 40-p = 65
p + 40 = 65
p = 25 (number of pigs)
.
Number of goats:
40-p = 40-25 = 15 (number of goats)